Problem 2 Salaries for new assistant professors at State College are normally distributed with a mean of $80000 and a standard deviation of $7000 We randomly survey 20 teachers from that district, let X be the salary of a randomly selected teacher at State, and let X be the average salary of 20 randomly selected teachers Expand All c. d3%) Find the probablity that ㄺ individual teacher earns more than $65,000. X> 65000) d. (3) Fied the probablity that the average salary for the sample is more than $65,000. PLX > 65000) e. 3%) Find the 90th percentle for the average teachers salary for samples of 20 teachers If PR < a) = 09 then
Answer #1

a)

X ~N(80000 , 7000)

b)Xbar ~ N(80000 , 1565.248)

c)

P(X>65000)=P(Z>(65000-80000)/7000)=P(Z>-2.14)=0.9838

d)

P(X>65000)=P(Z>(65000-80000)/1565.248)=P(Z>-9.58)=1.0000

e)for 90th percentile ; critical z =1.28

hence corresponding value a=65000+1.28*1565.248=67003.52